OpenStudy (anonymous):

13=6 mod A 17=8 mod B 3=1 mod C 11=5 mod D 9=4 mod E 15=7 mod F

OpenStudy (anonymous):

A = 1/2 B = 1/2 C = 1/2 D = 1/2 E = 1/2 F= 1/2

OpenStudy (anonymous):

If 'mod' means modulus: \[13=6|A|\]\[|A|=\frac{13}{6}\]\[A=\pm\frac{13}{6}\] If 'mod' means modulo: I have no idea what to do.

OpenStudy (anonymous):

I am trying to get the answers to solve this GPS coordinate (part of a puzzle) N 36 43. ABC W 076 20. DEF

OpenStudy (anonymous):

It does mean modulus

OpenStudy (amistre64):

13=6 mod A 1 2 3 4 5 6 7 8 9 10111213 14 mod 7 if i see it right

OpenStudy (amistre64):

17=8 mod B 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 B = 9 ??

OpenStudy (anonymous):

36 43.79C

OpenStudy (amistre64):

well if its right, then we can figure out the rest :) C 12 <-: 3 = 1 mod 2 34 11=5 mod D 1 2 3 4 5 6 <-- 11 = 5 mod 6 7 8 9 10 11 9=4 mod E 1 2 3 4 5 <- 9=4 mod 5 6 7 8 9 15=7 mod F 1 2 3 4 5 6 7 8 <- 15=7 mod 8 9 10 11 12 13 14 15

OpenStudy (anonymous):

amistre64.... you rock. Thanks I plugged in the numbers checked the coordinates on Geochecker.com and they were correct. I really appreciate the help!